OpenStudy (anonymous):

-0.4(n+4)=0.1(n-1.5) Whats n?

OpenStudy (anonymous):

-0.4n-0.16=0.1n-0.15 -0.4n-0.1n=-0.15+0.16 -0.5n=0.1 n= -0.1/0.5 =-1/5 ==-0.2

OpenStudy (anonymous):

you was wrong